zarkam21:

What is the power of a lightbulb that has a current of 0.480 A at 120.0 V? 120/0.480=250

Vocaloid:

P = IV so you would multiply not divide

zarkam21:

120.0*0.480=57.6

Vocaloid:

good, units of power are W (watts) as usual

zarkam21:

What is the voltage drop across a 40.0 Ω resistor that has a 0.350 A current?

zarkam21:

40/0.350=114.3

Vocaloid:

V = IR so once more you would multiply current and resistance not divide

zarkam21:

40*0.350=14

Vocaloid:

good, units of potential drop is V as usual